Business Costs Product or Period Cost Variable or Fixed Cost Per Unit Cost (Variable)
Annual Cost (Fixed)
Artist - contract Product Fixed $10,000
Artist - design fee Product Fixed $3,600
Computer/Printer 90% Product Fixed $5,400
Computer/Printer 10% Period Fixed $600
Depreciation - Computer/Printer 90% Product Fixed $1,800
Depreciation - Computer/Printer 10% Period Fixed $200
Depreciation - Heat Press Machine Product Fixed $1,500
Heat Press Machine Product Fixed $4,500
Liability Insurance Period Fixed $3,600
Mall Rent (Manufacutring) 90% Product Fixed $27,000
Mall Rent (Non-Manufacturing) 10% Period Fixed $3,000
Owners (shared) Period Fixed $12,000
Parties Period Fixed $4,000
Storage Unit (Manufacturing) Product Fixed $1,500
Inkjet Cartridges Product Variable $0.10
Inkjet Cartridges Period Variable $0.02
Laser Paper Period Variable $0.02
Student Workers (3) Product Variable $0.80
Student Workers (3) Period Variable $0.40
T-shirts Product Variable $3.75
Transfer Paper Product Variable $0.40
Wrapping Paper/Box Period Variable $0.20
Fixed (a) = $68,200.00 (included list of fixed business costs)
Variable (b) = $5.69 (included list of variable business costs)
Y =68200+5.69X

Questions are- Please answer with how on however many you can!

1. Assume that you make and sell 7,800 t-shirts in the first year. Use your cost formula to calculate your first year’s total cost. If you sell these t-shirts at $15 each, how much would net profit be in the first y e a r ? Year ending 12/31/2020.

Now you have developed your cost estimates, let’s do some evaluations on this proposed business. a. Continue to assume that 7,800 t-shirts will be made and sold in the first year. What is your product cost per unit under absorption costing? What is your product cost per unit under variable costing?

b. Based on the estimated sales level of 7,800 t-shirts for the first year, prepare your company’s (forecasted) income statement for the year ended on 12/31/2020 using both (1) the traditional format based on the absorption costing and (2) the contribution format based on the variable costing.

c. Calculate contribution margin per T-shirt and contribution margin ratio. d. Calculate how many T-shirts you need to sell in order to break-even. Calculate how much sales in dollars you need to make in order to break-even. (Use break-even f o r m u l a s .)

e. Calculate how many T-shirts you need to sell in order to make $10,000 target profit for the year.

f. Continue to assume that 7,800 T-shirts will be made and sold during the first year. Calculate your (1) margin of safety and (2) degree of operating leverage (DOL) for your business. What do these figures tell you about how risky your business is?
Answer #1

1. Cost formula: 68200+5.69X

7,800 t-shirts for the year ending 12/31/2020 -

Total cost: 68,200+5.69(7800) = 112,582$

Total Income: 15(7800) = 117,000$

Net Profit: 117,000 - 112,582 = 4,418$

a. Under absorption costing, the product cost absorbs all the cost of manufacturing a product, including direct material, direct labour, variable overheads, fixed overheads.

Cost under absorption costing:

Fixed Cost p.u.: 68,200/7800units = 8.74$ p.u.

Variable cost: 5.69$ p.u.

Total cost under absorption cost: 14.43$ p.u.

Under variable costing, the product cost includes only the variable cost of manufacturing the product including direct material, direct labour and variable overheads.

Total cost under variable costing: 5.69$ p.u.

b. 1. Income statement for the period ended 31/12/2020 under traditional format based on absorption method

Sales (7800*15$) 117,000$

Cost of Goods Sold (7800*5.69$) (44,382)$

(68,200)$

Gross Profit 4,418$  

2. Income statement for the period ended 31/12/2020 under contribution method

Sales (7800*15$) 117,000$

Variable cost (7800*5.69$) (44,382)$

Contribution 72,618$

Fixed cost (68,200)$

Net Income 4,418$

C. Contribution margin per t-shirt:

Sales p.u. 15$

Variable cost p.u   5.69$

Contribution p.u. 9.31$

Contribution margin ratio= Contribution / Sales

= 9.31/15 = 0.62

D. Break even point is the point of sale, where complete fixed cost is recovered by the organisation with no profit no loss scenario.

Break even point = Total Fixed cost / (Sales - Variable cost)

= 68,200 / (15$ - 5.69$)

= 7,326 Units

E. Target 10,000$ profit

Contribution    9.31$X

Fixed Cost (68,200)

Income 10,000

9.31X - 68,200 =10,000

9.31X = 78,200

X= 8,400 units

F. 1. Margin of Safety is the excess level of sales over and above break even point.

Margin of safety = (Current sales - break even point) / Current sales * 100

= (7800 - 7326) / 7800 *100

= 6.07%